Suppose a college decide to raise the tuition fee from Rs.3, 000 per month to Rs.4, 000 of MBA level students. The number of students getting admission is 500. The price elasticity of demand for the students is 0.2.

Find out the number of students getting admission after increase in tuition fee.

Find out the total revenue before and after the increase in tuition fee.

Explain whether the increase in tuition fee is wise decision or not on the basis of total revenue.


1
Expert's answer
2021-07-22T10:23:16-0400

a) Given:

Ed=0.2, Q1=500, P1=Rs. 3,000, P2=Rs. 4,000

We know that,

"E_d=\frac{∆Q}{∆P}×\frac{P}{Q};\space where\space ∆Q=Q_2−Q_1 \space\space and\space ∆P=P_2−P_1"


substituting the values given in the formula,

"E_d=−\frac{∆Q}{∆P}×\frac{P_1}{Q_1}"


"0.2=−\frac{∆Q}{(4,000−3,000)}×\frac{3,000}{500}"


"0.2=−\frac{∆Q}{1,000}×6\\∆Q=−\frac{0.2×1,000}{6}=−33.33"


Since,

"∆Q=Q_2-Q_1\\\Delta Q_2=\Delta Q+Q_1\\Q_2−33.33+500\\Q_2=466.67"


Therefore, the number of students getting admission after increase in tuition fee will be 466.67.


b) Total revenue (TR) is the product of price and quantity.

"=3,000×500\\=15,00,000"

Total Revenue before tuition fee rise will be Rs. 15 Lakh


After price rise, fees is 4000 and students are 466.67

TR after fee rise

"=4,000×466.67\\=18,66,680"

 

Total Revenue after tuition fee rise will be Rs. 18,66,680.


c)

18,66,680>15,00,000

The total revenue after fee rise is greater than the total revenue before fee rise.

Therefore, increasing the tuition fee will be a wise decision.